In a competitive market, the market demand is Qd = 60 - 6P and the market supply is Qs = 4P.

  1. Find the Equilibrium Price (Ep)
  2. Determine the Quantity Supplied (Qs)
  3. Determine the Quantity Demanded (Qd)
Answer #1

1) Put Qd= Qs

60-6P= 4P

60= 10P

P= 6

2)Qs= 4P = 24

3) Qd= 60-6P= 60-36=24

If you have any doubt feel free to ask.